OpenStudy (anonymous):

differentiate 2x^3+10/9x

OpenStudy (anonymous):

6x^2+10/9

OpenStudy (anonymous):

\[\huge (2x)^3 + \frac {10} { 9x } \\ \huge => 8x^3 + \frac { 10} {9x} \]

OpenStudy (anonymous):

@Rohangrr I believe it's 2(x^3) not the entire quantity cubed. But I could be wrong.

OpenStudy (anonymous):

it should be 6x^2 but I cant get the last part right

OpenStudy (anonymous):

It should just be 10/9

OpenStudy (anonymous):

would it be 90x^-1/2

OpenStudy (anonymous):

I tired and it said the answer was not correct

OpenStudy (anonymous):

2^3 = 2*2*2 = 8 Hence @Rohangrr is correct

OpenStudy (anonymous):

Its a differentiate problem so the power would drop one and be multiplied to the base

OpenStudy (anonymous):

@MathTHEORY It is 2(x^3) not (2x)^3 And when deriving (10/9)x=10/9

OpenStudy (anonymous):

So the differentiation is: \[6x^{2}-\frac{ 10 }{ 9x^{2} }\]

OpenStudy (anonymous):

I tried that earlier, it didnt work

OpenStudy (anonymous):

@marlrich The differentiation of 10/9x is: